OpenStudy (anonymous):

how much is a nanolight-year and how would i convert miles to nanolight years?

OpenStudy (anonymous):

I've not actually seen that before, but a light year is a measure of distance, so I'd guess a nano light year is as well. 1 light year = \(5.879\times10^{12}\)miles and typically nano indicates \(10^{-9}\), so I'd guess 1 light year = \(10^{-9}\) nano light year. So I think you can say \[5.870\times10^{12}~miles=10^{-9}~nano~light~years\] or, dividing by \(10^{-9}\) \[5.870\times10^{21}~miles=1~nano~light~year\]

OpenStudy (mathstudent55):

@peachpi A nano light year is a very small length (one billionth) compared to a light year, so 10^9 nano light years = 1 light year.

OpenStudy (mathstudent55):

\(10^{-9}\) light year = 1 nano light year or 1 light year = \(10^9\) nano light years

OpenStudy (anonymous):

ha. you are correct. I reversed them

OpenStudy (anonymous):

so can use a a conversion factor 1 mile = 5.87 x 10³ nano light year?

OpenStudy (mathstudent55):

\(\large 1~ nanolight~year \times \dfrac{1~light~year}{10^9~nanolight~year} \times \dfrac{5.879 \times 10^{12}~miles}{1~light~year}\)

OpenStudy (mathstudent55):

Instead of figuring out a conversion factor from miles to nanolight years, I prefer going through the units whose conversions I know.

OpenStudy (anonymous):

my assignment is to convert a random value for miles, convert it to nanolight years to find nanolight years per minute

OpenStudy (mathstudent55):

Then use the number of nanolight years you are given instead of 1 at the beginning of my conversion above.

OpenStudy (anonymous):

can you give me an example. im kind of confused

OpenStudy (mathstudent55):

Wait. To convert miles to nanolight-years do this:

OpenStudy (mathstudent55):

Let's say you want to convert 1000 miles to nanolight years. Start with 1000 miles: \(1000~miles\) Now apply the conversion factors as fractions: \(1000~miles \times \dfrac{1~light~year}{5.879 \times 10^{12}~miles} \times \dfrac{10^9 ~nanolight~year}{1~light ~year} \)

OpenStudy (mathstudent55):

Multiply all the numerators together and divide by all the denominators. \(= \dfrac{1000 \times 1 \times 10^9}{5.879 \times 10^{12} \times 1} ~nanolight~years\) \(= \dfrac{10^{12}}{5.879 \times 10^{12}} ~nanolight~years\) \(= 0.17 ~nanolight~years\)